# Utilizing the Janus MoSSe surface polarization property in complementary metal–oxide–semiconductor field effect transistor design

### v1

@@ Abstract
Janus transition metal dichalcogenides (JTMDs) have attracted much attention because of their outstanding electronic and optical properties. The additional out-of-plane dipole in JTMDs can form n- and p-like Ohmic contacts, and this may be used in device applications such as pin diodes and photovoltaic cells. In this study, we exploit this property to design n- and p-type metal–oxide–semiconductor field effect transistors (MOSFETs). First, we use density-functional theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe. The intrinsic dipole of MoSSe causes band bending at both the metal/MoSSe and MoSSe/metal interfaces, resulting in electron and hole accumulation to form n- and p-type Ohmic contact regions. We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, on the basis of which we design complementary MOSFETs. Our results demonstrate that JTMDs can be used to make n- and p-MOSFETs in the same layer without the need for any extra doping.
@@ I Introduction
In the field of two-dimensional (2D) semiconducting materials, research on transition metal dichalcogenides (TMDs) has been flourishing in recent years owing to their remarkable properties, such as relatively high carrier mobilities,Chhowalla, Jena, and Zhang (2016a); Chen et al. (2022a); Radisavljevic et al. (2011) suitable bandgap energies,Ye, Zhang, and Yap (2017); Chaves et al. (2020) and an atomic thickness that is modifiable.Serna et al. (2016); Zhang et al. (2013); Orofeo et al. (2014) These properties make 2D TMDs suitable for various device applicationsLiu et al. (2021); Cho, Park, and Kwon (2015); Ying-Chuan_KMC; Xiong et al. (2021); 2D_Photo, especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs). Liu et al. (2021); Chhowalla, Jena, and Zhang (2016b); 2D_FET_Pin-Fang These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials. For example, MoS and WS are two commonly selected materials for n-type channels, and -MoTe and WSe are both good candidates for p-type channels.Radisavljevic et al. (2011); Pezeshki et al. (2016); Ovchinnikov et al. (2014); Cho, Park, and Kwon (2015); Zhang et al. (2021)
@@
It is well known that the contact characteristics of a FET device will strongly affect its – characteristics and performance. For low contact resistance, Ohmic contacts are usually required at the semiconductor–metal interface to enhance the electrical performance of FETs while maintaining a linear – curve.Sze, Li, and Ng (2021); Allain et al. (2015); Zheng et al. (2021) Unfortunately, Schottky barriers (SBs) are usually formed for most TMD/metal contacts.Allain et al. (2015); Kang et al. (2014); Ouyang, Xiong, and Jing (2018); Tang et al. (2019); Zheng et al. (2021) For an FET with SB contacts, carriers must overcome the energy barrier to transport from metal into the TMD channel region, impairing the device function by lowering the injection efficiency.Tung (2014) A typical technique used to reduce contact resistance in modern semiconductor technology is substitutional doping.Younas, Zhou, and Hinkle (2023) However, there are better ways to apply such a doping method to 2D materials, since their characteristics will be heavily altered after doping. Additionally, the absence of dangling bonds on the surface of a 2D material makes solid interface connections with metal challenging, resulting in larger contact resistance.Allain et al. (2015) Hence, most of the effort required when engineering TMD/metal contacts is associated with finding a suitable metal to adjust the work function. For n-FETs, a material with a smaller work function is needed, whereas a metal with a high work function is often selected for p-FETs. Several studies have shown that TMD/metal heterostructures such as MoS/Cu, WS/Cu, and WSe/Ti could provide n-type SB contacts,Kang et al. (2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe/Pd and WSe/Au.Kang et al. (2014); Kong et al. (2020) However, only a specific TMD/metal combination is acceptable for Ohmic contact formation, making TMD-based FET contact design more challenging.Zhao and Schwingenschlögl (2022)
@@
To combine n- and p-type FETs into a complementary metal–oxide–semiconductor FET (CMOSFET) device, horizontally combing two different TMD channels to form a CMOSFET is a common approach.Liu et al. (2014); Das, Dubey, and Roelofs (2014); Cho, Park, and Kwon (2015); Xiong et al. (2021) With regard to fabrication, however, this CMOSFET design with two different materials has the disadvantage that the two materials need to be transferred onto a single substrate, with the risk of introducing defects or contamination.Ullah et al. (2021) Furthermore, two contact metals need to be considered separately, resulting in higher processing complexity and a more time-consuming fabrication procedure. Recently, ambipolar TMDs such as MoTeLin et al. (2014) and ReSeLee et al. (2019) have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
@@
The use of a novel type of materials, namely, Janus transition metal dichalcogenides (JTMDs), MXY, where M transition metal and X, Y S, Se, or Te, with X Y, is another possible way to fabricate a CMOSFET using a single material. In recent years, several JTMDs have been successfully synthesized via different techniques. Monolayer MoSSe, for instance, was synthesized utilizing chemical vapor deposition combined with plasma strippingLu et al. (2017) or a selective epitaxy atomic replacement process,Trivedi et al. (2020) while Lin et al. Lin et al. (2020) prepared Janus WSSe by pulsed laser deposition. Previous theoretical studies of these JTMDs have provided unmistakable proof that their phonon spectra are stable and devoid of imaginary frequencies, verifying their fabrication stability.Petrić et al. (2021); Chaurasiya, Gupta, and Dixit (2019) Compared with TMDs, JTMDs have two different chalcogen-atom layers, which create an asymmetric sandwich configuration that provides an out-of-plane electric field. This out-of-plane field has the characteristic of a dipole, which enables potential applications in piezoelectric, spintronics, catalysts, and gas sensing devices.Zhang et al. (2022, 2020); Kong et al. (2023) Also, simulation studies have shown that this intrinsic dipole feature of JTMDs can enable the formation of both n- and p-type contacts simultaneously without the need for additional doping when different sides are attached to graphene,Chen et al. (2022b); Palsgaard et al. (2018); Liu et al. (2020) and this can be incorporated into the design of pin-diodePalsgaard et al. (2018) or photovoltaicLiu et al. (2020) devices. Moreover, theoretical studies have shown that the SB at contact could be substantially reduced when JTMD MoSSe are attached to metals,Zhao and Schwingenschlögl (2022) owing to the intrinsic dipole. The above-mentioned JTMD/electrode contact systems may even attain Ohmic contact under certain conditions.Chen et al. (2022b); Zhao and Schwingenschlögl (2022) However, unlike optoelectronic devices, the exploitation of the advantages of n- and p-type Ohmic contacts in FET applications has not yet been explored. Furthermore, there has not been much work on utilizing this intrinsic dipole property to form FETs or CMOSFETs without doping.
@@
In this study, we propose a CMOSFET design concept utilizing the intrinsic dipole of a JTMD. The material MoSSe is selected for investigation, since it was the first JTMD and is the most commonly synthesized.Lu et al. (2017); Trivedi et al. (2020) The remainder of this paper is organized as follows. First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, bilayer MoSSe (2L-MoSSe)/graphene and MoSSe/Al (111) face and MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT). Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson–drift–diffusion solver (2D-DDCC) developed in our laboratory.Wu
@@ II.1 Heterostructure and layered MoSSe DFT simulations
The heterostructure calculations are performed with Quantum ESPRESSO,Giannozzi et al. (2009) an open-source DFT calculation software package employing the generalized gradient approximation (GGA) of the exchange-correlation functional and norm-conserving VanderbiltHamann, Schlüter, and Chiang (1979) pseudopotentials combined with DFT-D3 van der Waals correction.Grimme et al. (2010) An Monkhorst–Pack sampling is used. The geometry is refined until the maximal residual forces are less than 0.015 eV/Å, at which point MoSSe heterostructures are subjected to the dipole correction to negate the artificial electric field caused by their asymmetric structures. The vacuum region is set to 20 Å to prevent interactions between nearby supercells. For MoSSe contact simulation, we attach a 1 supercell of graphene to a two-layer AA stacking supercell of MoSSe, A supercell of six layers of (111) Al to a 1 supercell of MoSSe, and a 1 supercell of six layers of (111) Cu to a supercell of MoSSe. A simulation model is created for each metal’s energetically advantageous surface,Skriver and Rosengaard (1992) and four outward metal layers are fixed in the relaxation process. For layered MoSSe, we construct an AAA stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here. The only difference is that the mesh is set to in the MoSSe band structure calculation.
@@ II.2 Device simulation
The 2D-DDCC solver is used to examine the device performance. This solver was developed in our laboratory and has been used on many different kinds of devices. More detailed information can be found on our website.Wu Both the Poisson and drift–diffusion equations are solved self-consistently by the solver. The equations are as follows:
@@
where is the dielectric constant, is the electric potential, is the acceptor’s doping concentration, is the donor’s doping concentration, is the electron concentration, is the hole concentration, and are the current densities of electrons and holes, respectively, and and are the electron and hole diffusion coefficients, respectively. This solver serves as the basis for all of the device simulations discussed in this article, with the exception of the MoSSe channel region calculations.
III Results and discussion
To study the interactions between MoSSe and contacts, we construct the S and Se sides of the MoSSe heterostructure with graphene and two common electrode metals (namely, Al and Cu). Figure 1(a) shows the calculated heterostructure systems that can form low SB or Ohmic contacts. The top views are shown in Figs. 1(b) and 1(c), where the lattice mismatches are 1.11%, 0.44%, and 2.97% for the 2L-MoSSe/graphene, MoSSe/Al, and MoSSe/Cu systems, respectively. Figures 1(e)–1(j) show the projected band structures, and Figs. 1(e)–1(g) representing the S-face contact of MoSSe with the electrode material, and Figs. 1(h)–1(j) representing the Se-face attachment. As can be observed, if graphene contacts a two-layer MoSSe material on the Se side [Fig. 1(e)], then the Fermi level of the entire system is 0.1 eV lower than the observable conduction band of the MoSSe material, suggesting an n-type Schottky barrier. On the other hand, if graphene contacts a two-layer MoSSe material on the S side [Fig. 1(h)], then we see a p-type Ohmic contact. These results are consistent with the results of previous JTMD contact studies.Palsgaard et al. (2018); Chen et al. (2022b) According to our DFT calculation, monolayer MoSSe has a strong V/cm intrinsic dipole field in the out-of-plane direction, which may be the reason why this single material exhibits n- and p-type contact formation when the different sides of the MoSSe attach to graphene. Also, the Dirac cone of the graphene in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al. (2010)
We are also interested in metal electrode cases such as Al and Cu. For Al (111) face attachment to the S and Se faces of monolayer MoSSe [Figs. 1(f) and 1(i)], the MoSSe band intersects the Fermi level, indicating that Ohmic contact occurs between the metal and MoSSe states. The intersection can extend to the MoSSe conduction band when the S face attaches to Al, behaving as an n-type contact; p-type contact can be observed in the case of Se-face attachment, since the crossover point is mainly attributable to the valence band. Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111). In contrast to the 2L-MoSSe/graphene case, these metal contact projected band structures demonstrate the chemisorption characteristics due to significant hybridization in the band diagram. This chemisorption-like property imparts an the advantage with regard to tunneling barrier reduction at the contact region, thus improving device performance.Gong et al. (2010); Zhao and Schwingenschlögl (2022)
DFT calculations are also used to obtain the device simulation parameters and the electronic parameters MoSSe, such as bandgap, affinity, dielectric constant, and effective mass. Figure 2(a) shows the band diagram of the AAA MoSSe stacking with a relaxed lattice constant of 3.25 Å and an interlayer distance of 3.19 Å. All three layers exhibit an indirect bandgap with a valley in the conduction bands and a valley in the valence bands. The projected band structure of each MoSSe layer can be observed in Fig. 2(b). The yellow, blue, and red dots represents the projections of the top, middle, and bottom MoSSe layers, respectively, and are correlated with the colored arrows in Fig. 2(a). As the layers’ band components were separated, we observed that the bands were shifted by 0.43 eV as each MoSSe layer was added. This effect originates from the intrinsic dipole in the out-of-plane direction of MoSSe. Similar to the case of gallium nitride, we can represent this dipole effect by a spontaneous polarization term in the device simulations, which is illustrated by the two diagonal black arrows in Fig. 2(b). The term for MoSSe can be written as the product of the out-of-plane dielectric constant and the intrinsic electric field :
It is easy to obtain from the thickness of the MoSSe and the value of the out-of-plane dipole. However, a problem arises when attempting to obtain in DFT calculations, because the vacuum layer along the axis will also be included. The principle of the equivalent capacitance methodLaturia, de Put, and Vandenberghe (2018) is adopted here to separate the dielectric constants of MoSSe from those of the whole system. Carrier mobility is obtained using the Monte Carlo procedure described by Pai and Wu.Pai and Wu (2022) We design our device using MoSSe of 5 nm channel thickness to represent MoSSe multilayers, since this thickness ensures there is sufficient charge flowing in the channel region. All the parameters used in the 2D-DDCC simulation are given in Table 1. In our device simulation setting, the metal work function is set at the mid-gap of the MoSSe to showcase the possibility of n- and p-type contact formation even in the worst case metal selection scenario.
@@ III Results and discussion
| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO and AlO parameters are given inherently in 2D-DDCC.Wu |
@@
| 667 cm/(Vs) | Electron mobility | |
| 624 cm/(Vs) | Hole mobility | |
@@
| cm | Spontaneous polarization | |
| SiO dielectric constant | ||
| AlO dielectric constant | ||
| 5.8 eV | HfO bandgapWu, Alivov, and Negi (2008) | |
| HfO dielectric constantKumar, Kaushik, and Negi (2014) |

### v2

@@ Abstract
Janus transition metal dichalcogenides (JTMDs) have attracted much attention because of their outstanding electronic and optical properties. The additional out-of-plane dipole in JTMDs can form n- and p-like contacts, and this may be used in device applications such as pin diodes and photovoltaic cells. In this study, we exploit this property to design n- and p-type metal–oxide–semiconductor field effect transistors (MOSFETs). First, we use density-functional theory calculations to study the inherent dipole field strength in the trilayer JTMD MoSSe. The intrinsic dipole of MoSSe causes band bending at both the metal/SMoSe and SMoSe/metal interfaces, resulting in electron and hole accumulation to form p-type, n-type, and even Ohmic contact regions. We incorporate this property into a 2D finite-element-based Poisson–drift–diffusion solver to perform simulations, based on which we design complementary MOSFETs. Our results demonstrate that JTMDs can be used to make n- and p-MOSFETs in the same layer without the need for any extra doping.
@@ I Introduction
In the field of two-dimensional (2D) semiconducting materials, research on transition metal dichalcogenides (TMDs) has been flourishing in recent years owing to their remarkable properties, such as relatively high carrier mobilities,Chhowalla, Jena, and Zhang (2016a); Chen et al. (2022a); Radisavljevic et al. (2011) suitable bandgap energies,Ye, Zhang, and Yap (2017); Chaves et al. (2020) and an atomic thickness that is modifiable.Serna et al. (2016); Zhang et al. (2013); Orofeo et al. (2014) These properties make 2D TMDs suitable for various device applicationsLiu et al. (2021); Cho, Park, and Kwon (2015); Chen et al. (2023); Xiong et al. (2021); Zhao et al. (2021), especially for replacing traditional channel materials and overcoming the barrier of short-channel effects in smaller field effect transistors (FETs). Liu et al. (2021); Chhowalla, Jena, and Zhang (2016b); Chen, Chen, and Wu (2022) These TMD-based channel FETs have been studied for over a decade, and FETs have been successfully fabricated with TMD materials. For example, MoS2 and WS2 are two commonly selected materials for n-type channels, and -MoTe2 and WSe2 are both good candidates for p-type channels.Radisavljevic et al. (2011); Pezeshki et al. (2016); Ovchinnikov et al. (2014); Cho, Park, and Kwon (2015); Zhang et al. (2021)
@@
It is well known that the contact characteristics of a FET device will strongly affect its – characteristics and performance. For low contact resistance, Ohmic contacts are usually required at the semiconductor–metal interface to enhance the electrical performance of FETs while maintaining a linear – curve.Sze, Li, and Ng (2021); Allain et al. (2015); Zheng et al. (2021) Unfortunately, Schottky barriers (SBs) are usually formed for most TMD/metal contacts.Allain et al. (2015); Kang et al. (2014); Ouyang, Xiong, and Jing (2018); Tang et al. (2019); Zheng et al. (2021) For a FET with SB contacts, carriers must overcome the energy barrier to transport from metal into the TMD channel region, impairing the device function by lowering the injection efficiency.Tung (2014) A typical technique used to reduce contact resistance in modern semiconductor technology is substitutional doping.Younas, Zhou, and Hinkle (2023) However, there are better ways to apply such a doping method to 2D materials since their characteristics will be heavily altered after doping. Additionally, the absence of dangling bonds on the surface of a 2D material makes solid interface connections with metal challenging, resulting in larger contact resistance.Allain et al. (2015) Hence, most of the effort required when engineering TMD/metal contacts is associated with finding a suitable metal to adjust the work function. For n-FETs, a material with a smaller work function is needed, whereas a metal with a high work function is often selected for p-FETs. Several studies have shown that TMD/metal heterostructures such as MoS2/Cu, WS2/Cu, and WSe2/Ti could provide n-type SB contacts,Kang et al. (2014); Ouyang, Xiong, and Jing (2018) and p-type SB contacts can be formed with a few combinations such as WSe2/Pd and WSe2/Au.Kang et al. (2014); Kong et al. (2020) However, only a specific TMD/metal combination is acceptable for Ohmic contact formation, making TMD-based FET contact design more challenging.Zhao and Schwingenschlögl (2022)
@@
To combine n- and p-type FETs into a complementary metal–oxide–semiconductor FET (CMOSFET) device, horizontally combing two different TMD channels to form a CMOSFET is a common approach.Liu et al. (2014); Das, Dubey, and Roelofs (2014); Cho, Park, and Kwon (2015); Xiong et al. (2021) With regard to fabrication, however, this CMOSFET design with two different materials has the disadvantage that the two materials need to be transferred onto a single substrate, with the risk of introducing defects or contamination.Ullah et al. (2021) Furthermore, two contact metals need to be considered separately, resulting in higher processing complexity and a more time-consuming fabrication procedure. Recently, ambipolar TMDs such as MoTe2Lin et al. (2014) and ReSe2Lee et al. (2019) have been proposed to compose a single-channel material CMOSFET, which reduces the difficulty of device fabrication.
@@
The use of a novel type of materials, namely, Janus transition metal dichalcogenides (JTMDs), MXY, where M transition metal and X, Y S, Se, or Te, with X Y, is another possible way to fabricate a CMOSFET using a single material. In recent years, several JTMDs have been successfully synthesized via different techniques. Monolayer MoSSe, for instance, was synthesized utilizing chemical vapor deposition combined with plasma strippingLu et al. (2017) or a selective epitaxy atomic replacement process,Trivedi et al. (2020) while Lin et al. Lin et al. (2020) prepared Janus WSSe by pulsed laser deposition. Previous theoretical studies of these JTMDs have provided unmistakable proof that their phonon spectra are stable and devoid of imaginary frequencies, verifying their fabrication stability.Petrić et al. (2021); Chaurasiya, Gupta, and Dixit (2019) Compared with TMDs, JTMDs have two different chalcogen-atom layers, which create an asymmetric sandwich configuration that provides an out-of-plane electric field. This out-of-plane field has the characteristic of a dipole, which enables potential applications in piezoelectric, spintronics, catalysts, and gas sensing devices.Zhang et al. (2022, 2020); Kong et al. (2023) Also, simulation studies have shown that this intrinsic dipole feature of JTMDs can enable the formation of both n- and p-type contacts simultaneously without the need for additional doping when different sides are attached to graphene (Gr),Chen et al. (2022b); Palsgaard et al. (2018); Liu et al. (2020) and this can be incorporated into the design of pin-diodePalsgaard et al. (2018) or photovoltaicLiu et al. (2020) devices. Moreover, theoretical studies have shown that the SB at contact could be substantially reduced when JTMD MoSSe are attached to metals,Zhao and Schwingenschlögl (2022) owing to the intrinsic dipole. The above-mentioned JTMD/electrode contact systems may even attain Ohmic contact under certain conditions.Chen et al. (2022b); Zhao and Schwingenschlögl (2022) However, unlike optoelectronic devices, the exploitation of the advantages of n- and p-type Ohmic contacts in FET applications has not yet been explored. Furthermore, there has not been much work on utilizing this intrinsic dipole property to form FETs or CMOSFETs without doping.
@@
In this study, we propose a CMOSFET design concept utilizing the intrinsic dipole of a JTMD. The material MoSSe was selected for investigation since it was the first JTMD and is the most commonly synthesized.Lu et al. (2017); Trivedi et al. (2020) The remainder of this paper is organized as follows. First, we illustrate the JTMD dipole and contact characteristics for three heterostructures, namely, Gr/trilayer AA′A stacking MoSSe/Gr and A′A stacking bilayer MoSSe(2L-MoSSe)/Al (111) face and 2L-MoSSe/Cu (111) face structures, via first-principles density functional theory (DFT). Then, we demonstrate the JTMD CMOSFET device design concept with simulation results using a 2D finite-element-based Poisson–drift–diffusion solver (2D-DDCC) developed in our laboratory.Wu
@@ II.1 Heterostructure and layered MoSSe DFT simulations
The heterostructure calculations are performed with Quantum ESPRESSO,Giannozzi et al. (2009) an open-source DFT calculation software package employing the generalized gradient approximation (GGA) of the exchange-correlation functional and norm-conserving VanderbiltHamann, Schlüter, and Chiang (1979) pseudopotentials combined with DFT-D3 van der Waals correction.Grimme et al. (2010) An Monkhorst–Pack sampling is used. The geometry is refined until the maximal residual forces are less than 0.015 eV/Å. At this point, MoSSe heterostructures are subjected to dipole correction to negate the artificial electric field caused by their asymmetric structures. The vacuum region is set to 20 Å to prevent interactions between nearby supercells. For MoSSe contact simulation, we attach a 1 supercell of Gr to a 2L-MoSSe supercell. A supercell of six layers of (111) Al to a 1 2L-MoSSe supercell, and a 1 supercell of six layers of (111) Cu to a 2L-MoSSe supercell. A simulation model is created for each metal’s energetically advantageous surface,Skriver and Rosengaard (1992), and four outward metal layers are fixed in the relaxation process. For layered MoSSe, we construct an AA′A stacking of the MoSSe trilayer structure, with the above calculation strategy also being carried out here. The only difference is that the mesh is set to in the MoSSe band structure calculation.
@@ II.2 Device simulation
The 2D-DDCC solver is used to examine the device’s performance. This solver was developed in our laboratory and has been used on many different kinds of devices. More detailed information can be found on our website.Wu Both the Poisson and drift–diffusion equations are solved self-consistently by the solver. The equations are as follows:
@@
where is the dielectric constant, is the electric potential, is the acceptor’s doping concentration, is the donor’s doping concentration, is the electron concentration, is the hole concentration, and are the current densities of electrons and holes, respectively, and and are the electron and hole diffusion coefficients, respectively. This solver serves as the basis for all of the device simulations discussed in this article, except the MoSSe channel region calculations.
@@
| Parameter | Numerical value | Description111MoSSe parameters are all obtained by DFT. SiO2 and Al2O3 parameters are given inherently in 2D-DDCC.Wu |
@@
| 667 cm2/(Vs) | Electron mobility | |
| 624 cm2/(Vs) | Hole mobility | |
@@
| cm-2 | Spontaneous polarization | |
| SiO2 dielectric constant | ||
| Al2O3 dielectric constant | ||
| 5.8 eV | HfO2 bandgapWu, Alivov, and Negi (2008) | |
| HfO2 dielectric constantKumar, Kaushik, and Negi (2014) |
III Results and discussion
DFT calculations are also used to obtain the device simulation parameters and the electronic parameters MoSSe, such as bandgap, affinity, dielectric constant, and effective mass. Figure 1(a) shows the band diagram of the AA′A MoSSe stacking with a relaxed lattice constant of 3.25 Å and an interlayer distance of 3.19 Å. All three layers exhibit an indirect bandgap with a valley in the conduction bands and a valley in the valence bands. The projected band structure of each MoSSe layer can be observed in Fig. 1(b). The yellow, blue, and red dots represent the projections of the top, middle, and bottom MoSSe layers, respectively, and are correlated with the colored arrows in Fig. 1(a). As the layers’ band components were separated, we observed that the bands were shifted by 0.43 eV as each MoSSe layer was added. This effect originates from the intrinsic dipole in the out-of-plane direction of MoSSe. Similar to the case of gallium nitride, we can represent this dipole effect by a spontaneous polarization term in the device simulations, which is illustrated by the two diagonal black arrows in Fig. 1(b). The term for MoSSe can be written as the product of the out-of-plane dielectric constant and the intrinsic electric field :
It is easy to obtain from the thickness of the MoSSe and the value of the out-of-plane dipole. However, a problem arises when attempting to obtain in DFT calculations, because the vacuum layer along the axis will also be included. The principle of the equivalent capacitance methodLaturia, de Put, and Vandenberghe (2018) is adopted here to separate the dielectric constants of MoSSe from those of the whole system. Carrier mobility is obtained using the Monte Carlo procedure described by Pai and Wu.Pai and Wu (2022)
The aforementioned also influences the interactions between MoSSe and contacts. Initially, we construct the Gr/AA′A-MoSSe/Gr heterostructure, with the S and Se sides attached to the Gr electrodes, resulting in a lattice mismatch of 1.11% between MoSSe and Gr. In Figure 2(b), the projected band structures of the Gr/AA′A-MoSSe/Gr system are illustrated. Notably, when Gr contacts a three-layer MoSSe material on the Se side (depicted in red), the Fermi level of the entire system is observed to be 0.38 eV lower than the conduction band of the MoSSe material, indicating an n-type Schottky barrier. Conversely, when Gr contacts a three-layer MoSSe material on the S side (depicted in green), a p-type Ohmic contact is observed at the Gamma point. These findings align with previous studies on JTMD contacts.Palsgaard et al. (2018); Chen et al. (2022b) According to our DFT calculation, monolayer MoSSe exhibits a strong V/cm intrinsic dipole field in the out-of-plane direction. This characteristic may explain why this single material exhibits both n- and p-type contact formations when different sides of MoSSe attach to Gr. Furthermore, in Figure 2(c), the effect of the intrinsic dipole on the system is depicted; the pulls the Gr downward as the Se side contacts the Gr. Consequently, the Fermi level becomes higher than the Dirac cone, resulting in intrinsically n-doped Gr. Due to the tilting band edge caused by , the energy difference between the conduction band edge and the Fermi level is smaller than the valence band edge offset at the Se/Gr side, suggesting the formation of an n-region. On the other hand, the S-side attachment elevates the band energy of the Gr, with the valence band edge intersecting the Fermi level, indicating an intrinsic p-doped Gr and a p-contact region. Nevertheless, the Dirac cone of both Gr in the band structure retains its integrity, indicating the presence of physisorption characteristics between the two material systems.Gong et al. (2010)
We also simulate metal contact systems due to the interest in metal electrode cases, where common electrode materials such as Al and Cu are selected.Allain et al. (2015); Kang et al. (2014); Ouyang, Xiong, and Jing (2018); Tang et al. (2019); Zheng et al. (2021) Fig. 3(a) illustrates the simulated 2L-MoSSe/Metal heterostructure and the top views are shown in Figs. 3(b) and 3(c), where the lattice mismatches are 0.44%, and 2.97% for the 2L-MoSSe/Al and 2L-MoSSe/Cu systems. Figures 3(d)–3(g) show the projected band structures, and Figs. 3(d)–3(f) representing the Se-face contact of MoSSe with the electrode material, and Figs. 3(e)–3(g) representing the S-face attachment.
For Al (111) face attachment to the Se and S faces of bilayer MoSSe (2L-MoSSe) [Figs. 3(d) and 3(e)], the 2L-MoSSe band intersects the Fermi level, indicating that Ohmic contact occurs between the metal and 2L-MoSSe states. In contrast to the 2L-MoSSe/Gr cases, these metal contact projected band structures demonstrate the chemisorption characteristics due to significant hybridization in the band diagram. Different from typical TMDs, the internal electric field in JTMD plays a crucial role in augmenting the density of interface states. Specifically, in MoSSe, this electric field fosters the generation of electronic energy levels at the metal-semiconductor interface via chemisorption-like bonding, leading to an increased density of interface states between the two materials. Zhao and Schwingenschlögl (2022) These additional bonding states, resulting from in MoSSe, are delineated in density of states (DOS) plots for all systems, indicated by a dotted frame.Gong et al. (2010) Due to the additional bonding states caused by chemical adsorption, it becomes challenging to correlate the band edge of MoSSe closest to the metal layer.
We correlate the band edge distribution of the closest-to-metal layer of MoSSe (1L MoSSe) based on the direction of previously mentioned and the peak present at the Ev edge of MoSSe layers, and depict the corresponding band edge with DOS on the right side of Figs. 3(d) and (e). When the Se face contacts with Al [Fig. 3(d)], it can be observed that the Fermi level coincides with the outer MoSSe (2L MoSSe) at the valence band edge. Due to the upward direction of , the energy band of the inner layer of MoSSe (1L MoSSe) shifts towards lower energy. Supported by the peak at the Ev edge, it can be inferred that the position of the Fermi level is closer to the conduction band, leaning towards an n-type contact. Conversely, when the S face contacts with Al [Fig. 3(e)], with the downward direction of , the energy band of the 1L MoSSe shifts towards higher energy. A comparison reveals that the Fermi level is situated in the middle of the band, leaning towards a p-type contact. Although we correlate the n- or p-type contact in both systems, there are numerous bonding states present at the gap position of 1L MoSSe, indicating the Ohmic contact formation. In other words, both electrons and holes can freely flow through this low-tunneling barrier Ohmic contact, thereby enhancing device performance.Gong et al. (2010); Zhao and Schwingenschlögl (2022) Similar n- and p-type Ohmic contact formation results for different face attachments are also seen in the case of Cu (111) [Fig. 3(f) and (g)], indicating Cu can also bring Ohmic contact properties as 2L-MoSSe attches. The chemisorption-like property imparts the advantage concerning tunneling barrier reduction at the contact region, thus improving device performance. Zhao and Schwingenschlögl (2022); Gong et al. (2010)